OpenStudy (anonymous):

can't find the inverse even though tried 2-3 times. Can anybody help??

OpenStudy (anonymous):

The Matrix is 4 -1 1 2 2 3 5 -2 6

OpenStudy (anonymous):

There is a trick to finding matrix of 3x3.

OpenStudy (anonymous):

1) Find the determinate of the matrix.

OpenStudy (anonymous):

Can u show me??

OpenStudy (anonymous):

I have to solve this by adjoining it to an identity matrix

OpenStudy (anonymous):

2) The inverse = \[\frac{ 1 }{ \det(M) }\left[\begin{matrix}A & -B & C\\ -D & E & -F\\ G & -H & I\end{matrix}\right]\]

OpenStudy (anonymous):

Transpose the original matrix. \[\left[\begin{matrix}4 & 2 & 5 \\ -1 &2 &-2\\ 1 & 3 & 6\end{matrix}\right]\]

OpenStudy (anonymous):

A = \[\det \left[\begin{matrix}2 & -2 \\ 3 & 6\end{matrix}\right]\]

OpenStudy (anonymous):

B = \[\det \left[\begin{matrix}-1& -2 \\ 1 & 6\end{matrix}\right]\]

OpenStudy (anonymous):

C = \[\det \left[\begin{matrix}-1 & 2 \\ 1 & 3\end{matrix}\right]\]

OpenStudy (anonymous):

D = \[\det \left[\begin{matrix} 2& 5 \\ 3 & 6\end{matrix}\right]\]

OpenStudy (anonymous):

Do you see the pattern?

OpenStudy (anonymous):

Thanks for your help but this is not the method I have been taught. So, even though I might understand it, I cannot use it...

OpenStudy (anonymous):

You can solve the matrix by adjoining it with the identity or through this method. I did my linear algebra exam using this method.

OpenStudy (anonymous):

I hope this helps to check your answers. It's a fast way to check because you're only computing 2x2 determinants

OpenStudy (anonymous):

Thanks I'll keep that in mind...☺

OpenStudy (anonymous):

Can Anybody show solution by adjoining the identity matrix...???

OpenStudy (anonymous):

you can also solve it by row and column transformation method

OpenStudy (anonymous):

we have not been taught that method